Urgent Corporation had earnings per share of $4 last year, and it paid a $2 dividend.

Total retained earnings increased by $12 million during the year, and book value per

share at year-end were $40. Urgent Corporation has no preferred stock, and no new

common stock was issued during the year. If Argent’s year-end debt (which equals its

total liabilities) was $120 million, what was the company’s year-end debt/assets ratio?


1
Expert's answer
2021-12-23T16:19:53-0500

We will find the authorized capital

"Number of common stock = \frac{Retained earnings}{
(EPS-DPS)}=\frac{12}{4-2}=6"

"authorized capital=6\times40=240"

"EPS=\frac{net profit}{Number of common stock}"

"4=\\frac{net profit}{6}\n\n\u200b"

net profit=24

"Retained earnings=24+24-2\times6=42"

liabilities=assets=240+42=282


"debt/assets ratio=\frac{120}{282}=0.43" or 43%


43% of the company's assets are financed by debt